|
ID |
Date |
Author |
Subject |
Text |
 |
|
224
|
Tue Feb 22 02:02:11 2022 |
JM, CB, LV, PM, AK | Beam ON, Target ON Run 54 | 03:00
Beam ON ~7E6 pps
Target ON and open to the ring |
10x |
|
223
|
Tue Feb 22 00:32:35 2022 |
JM, CB, LV, PM, AK | Beam ON, Target ON Run 54 | 01:30
Beam ON ~6E6 pps
Target ON and open to the ring |
 |
|
222
|
Mon Feb 21 21:04:08 2022 |
SHB, JM, CB, OH (remotely) | Schottky RF view from the beam inside | |
|
|
221
|
Mon Feb 21 20:56:51 2022 |
SHB, JM, CB, OH (remotely) | DAQ connection for ion current | signals from ICT and CRYRADIO are now connected
to AIDA2 and AIDA5 respectively. ICT is set
to factor 200 and CRYRADIO to 40 for now. |
|
|
220
|
Mon Feb 21 18:03:58 2022 |
SHB, JM, CB, OH (remotely) | Info: signals on ion intensity for experiment | Currently 4 possibilities to have a look
at the beam, all of them are OK to check
whether there is beam or not, but some of |
|
|
219
|
Mon Feb 21 17:26:19 2022 |
JM, CB, SHB, OH (remotely) | Beam ON run 53 | 21/2/22 16:00
Call from main control room, beam
is ON, probably has been for several hours |
9x |
|
218
|
Mon Feb 21 12:10:47 2022 |
RSS, JG | Alpha background run50 | Beam OFF. Run started 20/2/22 18:42
System checck OK, except white
rabbit all fail |
|
|
217
|
Mon Feb 21 10:59:48 2022 |
RSS, JG | Alpha background run50 | Beam OFF. Run started 20/2/22 18:42
System checck OK, except white
rabbit all fail |
|
|
216
|
Mon Feb 21 09:07:37 2022 |
RSS, JG | Alpha background run50 | Beam OFF. Run started 20/2/22 18:42
System checck OK, except white
rabbit all fail |
|
|
215
|
Mon Feb 21 08:05:45 2022 |
RSS, JG | Alpha background run50 | Beam OFF. Run started 20/2/22 18:42
System checck OK, except white
rabbit all fail |
|
|
214
|
Sun Feb 20 22:04:21 2022 |
JM, CB | Alpha Background run 50 | Beam OFF. Run started 20/2/22 18:42
Fast comparator threshold HEC set
to 0xff for all fees and asics to combat |
|
|
213
|
Sat Feb 19 20:08:43 2022 |
JM, CB | Beam ON run 49 | Run started 19/2/22 21:05. Beam ON
Temperatures OK
System wide checks OK - white rabbit |
6x |
|
212
|
Sat Feb 19 14:51:58 2022 |
JM, CB | Burst high energy noise | Noticed bursts of high energy noise on
aida01 and aida02 during run 47 and 48.
Good events rate jumps from 10-20k |
    |
|
211
|
Fri Feb 18 20:31:33 2022 |
JM, CB | Alpha Background run 48 | Alpha background started with beam ON -
run47. Run started 16:49 18/2/22
(beam subsequently found to be OFF) |
|
|
210
|
Fri Feb 18 13:03:02 2022 |
RSS, ML, CB | Alpha background run46 | 14:00
Check Leakage current: (see attachment |
11x |
|
209
|
Fri Feb 18 11:07:52 2022 |
RSS, ML, CB | Alpha background run46 | 12:00
Check Leakage current: (see attachment |
11x |
|
208
|
Fri Feb 18 08:58:08 2022 |
RSS, ML, CB | Alpha background run46 | 10:00
Check Leakage current: (see attachment |
11x |
|
207
|
Thu Feb 17 18:19:16 2022 |
JM, CB, SD | Alpha Background run 46 | Both GSI vacuum group and edinburgh ion
gauges OFF
Fees power cycled. All system checks |
   |
|
206
|
Thu Feb 17 16:26:27 2022 |
JM, CB, SD | Gauges OFF, Pulser Movement tests | Edinburgh group ion gauge turned OFF ->
decrease in leakage current from 10.05 uA
to 6.5 uA |
7x |
|
205
|
Thu Feb 17 15:09:40 2022 |
JM,CB,SD | Alpha background run44 | 16:09
Check Leakage current: (see attachment |
8x |
|